Choose a generic formula for combination reaction: A) A+B=C B) AB=A+B C) A+BC=AC+B D) AB+CD=AD+CB OB OA
Answer #1

A) A+B= combination B) A B = A + B : Decomposition c) A + BC = AC + B : single replacement D AB + CD = AD + CB : Double repla